### v2

# PAFT: Prompt-Agnostic Fine-Tuning

Abstract
Fine-tuning large language models (LLMs) often causes overfitting to specific prompt wording, where minor phrasing variations drastically reduce performance. To address this, we propose Prompt-Agnostic Fine-Tuning (PAFT), a method that enhances robustness through dynamic prompt variation during training. PAFT first generates diverse synthetic prompts, then continuously samples from this set to construct training instances, forcing models to learn fundamental task principles rather than surface-level patterns. Across systematic evaluations using both supervised fine-tuning (SFT) and reinforcement learning fine-tuning (RLFT), PAFT demonstrates substantially improved prompt robustness, achieving 7% higher generalization accuracy on unseen prompts than standard methods. In addition to enhanced robustness, PAFT consistently yields superior overall performance on established benchmarks for question answering, mathematical reasoning, and tool use. Notably, models trained with PAFT attain 3.2× faster inference speeds due to reduced prompt sensitivity. Ablation studies further validate effectiveness of PAFT, while theoretical analysis reveals that PAFT can effectively enhance the cross-domain generalization ability of LLM. 00footnotetext: corresponding author.
PAFT: Prompt-Agnostic Fine-Tuning
Chenxing Wei†§, Mingwen Ou∘, Ying He#†, Yao Shu#≀, Fei Yu‡ †College of Computer Science and Software Engineering, Shenzhen University, China ∘Tsinghua Shenzhen International Graduate School, Tsinghua University, China §Guangdong Lab of AI and Digital Economy (SZ), China ≀Hong Kong University of Science and Technology (Guangzhou), China ‡School of Information Technology, Carleton University, Canada weichenxing2023@email.szu.edu.cn, yaoshu@hkust-gz.edu.cn
1 Introduction
Large language models (LLMs) have demonstrated remarkable success across diverse natural language processing (NLP) tasks Zhao et al. (2024); Xu et al. (2023). To further enhance the performance of LLMs on specific downstream tasks, supervised fine-tuning (SFT) Ouyang et al. (2022); Devlin et al. (2019) and reinforcement learning fine-tuning (RLFT) Wang et al. (2024); Wei et al. (2025) has emerged as a widely adopted strategy. These methods typically augment input data with task-specific instructions and construct dialogue datasets with expected outputs, enabling models to learn task-specific patterns. Empirical studies have shown that SFT and RLFT can substantially improve model performance on downstream tasks Raffel et al. (2023); Hu et al. (2023); Wei et al. (2022).
However, as shown in Figure 1, a critical limitation of current fine-tuning methods is their lack of prompt robustness, as further detailed in Sec. 3. Reliance on fixed instruction prompts Mishra et al. (2022); Chung et al. (2022) often leads to overfitting on specific prompts patterns Zhang et al. (2024); Kung and Peng (2023). Consequently, models become brittle: minor deviations between user and training prompts can significantly degrade inference performance Mialon et al. (2023); Raman et al. (2023). This brittleness manifests, for example, as substantial accuracy drops in QA tasks with altered prompt phrasing Wei et al. (2024), or as poor instruction following in chatbots and AI agents when commands deviate from those encountered during training Hong et al. (2024); Sahoo et al. (2025). Such sensitivity also raises fairness and reliability concerns in algorithmic comparisons Voronov et al. (2024). This vulnerability is particularly acute when users, unfamiliar with specific SFT prompt structures, provide highly divergent inputs, potentially causing fine-tuned models to perform near random guessing levels Polo et al. (2024). Notably, prompt robustness in SFT has received limited attention, with most existing work focusing on in-context learning and prompt tuning Shi et al. (2024); Ishibashi et al. (2023).
To address this critical gap, we introduce PAFT, a novel framework that dynamically adapts to diverse training prompts. To our knowledge, PAFT is the first systematic approach to improving prompt robustness in both SFT and RLFT, a vital but underexplored area. Unlike traditional methods prone to overfitting specific prompt patterns, PAFT enables models to grasp underlying task semantics, ensuring robust performance across varied human-written prompts. PAFT operates in two phases (Figure 2): first, constructing a diverse set of high-quality synthetic prompts that capture essential task semantics with linguistic variability (Sec. 4.1); second, employing dynamic fine-tuning by sampling from this curated set to expose the model to various formulations (Sec. 4.2). Extensive evaluations demonstrate that PAFT significantly boosts model robustness and generalization to diverse prompts, maintains state-of-the-art downstream performance, and can potentially improve inference speed while preserving training efficiency. These findings highlight PAFT as a promising direction for developing more robust, user-friendly language models.
Our key contributions are as follows: (a) We highlight that fine-tuning with fixed prompts results in poor generalization to unseen prompts and severe performance degradation (Sec. 3). (b) We propose PAFT, a novel framework incorporating candidate prompt construction and dynamic fine-tuning, to enhance the prompt robustness of fine-tuned models (Sec. 4). (c) We empirically demonstrate the consistent and robust performance of PAFT across diverse downstream tasks, fine-tuning algorithms, and varied test prompts, including those unseen during training (Sec. 5). (d) We provide theoretical evidence that PAFT effectively enhances the cross-domain generalization of LLMs (Sec. 6).
2 Related Work
Prompt Optimization.
Prompt engineering critically influences LLM performance, driving numerous prompt optimization approaches Chang et al. (2024); Li (2023); Diao et al. (2023); Sun et al. (2022). Notable methods include INSTINCT Lin et al. (2024), which leverages neural network bandits with LLM embeddings for search efficiency. ZOPO Hu et al. (2024), which employs localized search strategies. BATprompt Shi et al. (2024), which integrates robustness through natural language perturbations. While these approaches excel at identifying single high-performance prompts, models fine-tuned on such prompts remain vulnerable to prompt variations. Our work, in contrast, addresses this limitation by simultaneously enhancing prompt robustness and optimizing performance across the entire prompt space rather than focusing on isolated optimal prompts.
Fine-tuning (FT).
SFT and RLFT constitute the main paradigms for adapting LLMs to downstream tasks, prized for their efficiency. These approaches split into two categories: soft prompt tuning, which optimizes continuous input vectors while preserving base model parameters Kong et al. (2025); Wu et al. (2024); Hu et al. (2024); Lin et al. (2024); Li and Liang (2021); Liu et al. (2022), and full/parameter efficient fine-tuning (PEFT) Shu et al. (2024); Ouyang et al. (2022); Liu et al. (2021); Lester et al. (2021). Among PEFT techniques, Low-Rank Adaptation (LoRA) Hu et al. (2022) predominates by freezing pre-trained weights while introducing trainable low-rank matrices, with recent variants enhancing generalization and reducing overfitting Chen et al. (2023); Si et al. (2024); Wei et al. (2024). Instruction tuning Sanh et al. (2022) further improves ability of model to follow diverse task-specific instructions. However, existing methods—particularly soft prompt tuning—still exhibit limited prompt robustness, leaving models vulnerable to prompt variations. Our work addresses this critical limitation while maintaining computational efficiency.
3 Preliminaries
To systematically study the impact of prompt variations on fine-tuned models, we conducted comprehensive experiments across multiple downstream tasks using LLaMA3-8B Meta (2024) with LoRA fine-tuning. We constructed a comprehensive set of over 450 prompts (both human-written and LLM-generated), covering a wide range of language styles, task-specific instructions, and formatting variations. Figure 3 presents a statistical analysis of the accuracy distribution for both the base and SFT models across these prompts, revealing a key finding: the formulation of the prompt dramatically influences the performance of the model regardless of the type of task, with only 10% of the prompts producing near-optimal results. Minor prompt modifications (e.g., rephrasing, punctuation, reordering) induce substantial fluctuations.
For example, the addition of "Question" improves accuracy by 20% (Figure 1). This sensitivity highlights the fragility of current fine-tuning methods and their strong dependence on specific prompt formulations. These findings align with prior work He et al. (2024); Voronov et al. (2024); Salinas and Morstatter (2024); Min et al. (2022); Gao et al. (2021b). This widespread sensitivity demonstrates a fundamental limitation in current fine-tuning approaches, extending findings from previous research across diverse task domains. Based on these insights, we propose PAFT, which decouples model performance from specific prompt formulations, ensuring consistent results across prompt variations and enhancing practical applicability in real-world scenarios.
4 The PAFT Framework
In this section, we introduce PAFT in detail. As shown in Figure 2, the PAFT framework consists of two key stages: candidate prompt construction (Sec. 4.1) and dynamic fine-tuning (Sec. 4.2).
4.1 Candidate Prompt Construction
To ensure the robustness and effectiveness of PAFT across diverse prompts, we design a comprehensive prompt construction framework that aims to generate diverse and meaningful candidate prompts efficiently, enabling the model to generalize across different prompt formats. Our approach leverages the powerful generative capabilities of LLMs Kohl et al. (2024) and comprises three key phases.
Diverse LLM Ensemble. We employ 10 mainstream LLMs with varied generation capabilities OpenAI et al. (2024); Bai et al. (2023); Ouyang et al. (2022) to capture the inherent variability in task interpretation stemming from differences in pre-training data, architectures, and optimization objectives Minaee et al. (2024); Zhao et al. (2024). This diversity ensures comprehensive coverage of prompt formulations across linguistic styles and instructional approaches, effectively mitigating single-model generation biases.
Dual Prompting Strategy. We combine few-shot and zero-shot techniques to balance quality and diversity. Few-shot prompting leverages in-context learning with curated human examples to generate task-aligned, semantically coherent prompts. Zero-shot prompting encourages diverse linguistic styles and structural variations without explicit examples. By generating 20 prompts with each strategy, we create a comprehensive set spanning high-quality and varied formulations, exposing the model to realistic prompt quality distributions and enhancing robustness to real-world scenarios. See Appendix D.1 for details.
Rigorous Evaluation Design. We randomly partition generated prompts into training and test sets (8:1 ratio), ensuring completely distinct prompts in each set. This approach exposes the model to diverse prompt styles during training while providing a robust testbed for assessing generalization to novel formulations, see Appendix D for more details of prompt sets. By evaluating on entirely unseen prompts, we confirm that performance improvements reflect genuine ability to handle diverse prompt formulations rather than overfitting to specific patterns. This framework ensures PAFT learns task semantics independently of prompt phrasing, enabling effective generalization across real-world scenarios while providing a scalable, cost-effective solution for improving prompt robustness.
4.2 Dynamic Fine-Tuning
Dynamic Fine-Tuning Algorithm. Our PAFT framework enhances the robustness of LLMs through systematic prompt diversification. As shown in Algorithm 1, each training epoch randomly samples a prompt from synthetic candidates (line 4), exposing the model to varied linguistic styles. For each data point (line 6), the selected prompt is reused for consecutive steps (lines 7-9), constructing inputs via (line 7) and updating parameters using gradient-based optimization like SGD (Sra et al., 2011) or AdamW (Loshchilov and Hutter, 2019) (line 8). After steps, a new prompt is sampled (lines 10-11), ensuring multiple prompt exposures per epoch. Each epoch initializes with final parameters from the previous one: (line 12), maintaining learning continuity until final parameters are achieved after epochs (line 16).
Benefits of Dynamic Fine-Tuning. Dynamic fine-tuning in PAFT significantly enhances LLM robustness and generalization by exposing the model to diverse prompts during training. This approach mitigates overfitting to fixed prompts, fostering the learning of more generalizable representations less sensitive to specific formulations. Consequently, PAFT achieves consistent performance across varied and unseen prompts, crucial for real-world applications with diverse user input. By reducing reliance on manual prompt engineering, dynamic fine-tuning offers an efficient and scalable solution for improving LLM adaptability.
| Methods | Hellaswag | PIQA | Winogrande | RACE-mid | RACE-high | Average | ||||||||||||
| Metric | Mean | Std | Top | Mean | Std | Top | Mean | Std | Top | Mean | Std | Top | Mean | Std | Top | Mean | Std | Top |
| Base Model | 47.36 | 9.78 | 0% | 74.68 | 6.24 | 0% | 45.15 | 11.78 | 0% | 71.39 | 7.33 | 0% | 67.62 | 6.78 | 0% | 61.24 | 8.38 | 0% |
| User | 92.35 | 2.78 | 0% | 77.87 | 2.36 | 0% | 78.16 | 7.97 | 0% | 79.88 | 6.32 | 22% | 81.05 | 4.45 | 4% | 81.86 | 4.78 | 5% |
| TopAccuracy | 91.27 | 2.79 | 86% | 75.96 | 3.89 | 0% | 66.77 | 3.94 | 0% | 84.81 | 4.06 | 59% | 82.45 | 3.26 | 14% | 80.25 | 3.63 | 32% |
| BATprompt | 90.30 | 1.79 | 78% | 83.41 | 1.74 | 16% | 69.01 | 4.45 | 0% | 83.92 | 5.38 | 65% | 81.33 | 4.21 | 12% | 81.56 | 3.51 | 34% |
| ZOPO | 92.46 | 2.43 | 86% | 83.52 | 2.23 | 27% | 74.75 | 3.81 | 0% | 83.50 | 5.05 | 51% | 82.36 | 4.53 | 35% | 83.32 | 3.61 | 40% |
| PAFT | 93.83 | 0.70 | 100% | 89.33 | 0.63 | 100% | 82.09 | 0.81 | 100% | 87.26 | 2.23 | 94% | 85.17 | 1.71 | 73% | 87.57 | 1.57 | 94% |
| Improv. | +1.37 | -1.09 | 14% | +5.81 | -1.11 | 73% | +3.93 | -3.00 | 100% | +2.45 | -1.83 | 29% | +2.72 | -1.55 | 38% | +4.25 | -1.94 | 54% |
| Method | HumanEval | Xstory_cloze | Geometry3k | T-Eval | GSM8K |
| Base | 41.31 ( 10.36) | 48.23 ( 8.36) | 32.17 ( 15.36) | 58.97 ( 14.03) | 74.36 ( 21.37) |
| SFT | 49.63 ( 4.31) | 54.77 ( 4.79) | 37.94 ( 6.17) | 70.37 ( 8.14) | 81.47 ( 13.24) |
| PAFT | 54.24 ( 1.36) | 60.27 ( 0.73) | 40.19 ( 1.27) | 73.17 ( 3.27) | 85.71 ( 5.93) |
| Improv. | +4.61 (-2.95) | +5.50 (-4.06) | +2.25 (-4.90) | +2.80 (-4.87) | +4.24 (-7.31) |
| Inference time/h | Hellaswag | PIQA | Winogrande | RACE | Average |
| Base Model | 3.97 | 1.35 | 1.72 | 6.24 | 3.32 |
| User | 6.52 | 0.98 | 3.27 | 8.23 | 4.75 |
| TopAccuracy | 5.75 | 1.13 | 2.76 | 7.56 | 4.30 |
| BATprompt | 4.57 | 1.57 | 3.14 | 7.98 | 4.32 |
| ZOPO | 5.12 | 0.87 | 3.23 | 8.28 | 4.38 |
| PAFT | 1.19 | 0.39 | 0.45 | 2.08 | 1.02 |
| Improv. | 3.3 | 2.23 | 3.82 | 3.00 | 3.25 |
5 Empirical Results
We evaluate our PAFT framework through comprehensive experiments. Sec. 5.1 describes datasets and experimental setup, Sec. 5.2 analyzes key findings, and Sec. 5.3 presents ablation studies examining critical framework components.
5.1 Datasets and Setup
Benchmark Selection. As a pioneering work addressing prompt robustness in LLMs through training, we conduct experiments across a diverse set of tasks and benchmarks to ensure a comprehensive evaluation. Our methods involve Supervised Fine-Tuning (SFT) and a reinforcement learning approach, GRPO. For our SFT experiments, we selected benchmarks to cover a wide range of capabilities. Specifically, we use HellaSwag Zellers et al. (2019) for knowledge understanding, WinoGrande Sakaguchi et al. (2019) for language understanding, and RACE Lai et al. (2017) for reading reasoning capabilities. For grounding and abstractive summarization, we employ PIQA Bisk et al. (2019). To evaluate more specialized skills, we utilize HumanEval Chen et al. (2021) for coding, T-Eval Chen et al. (2024) for tool use, and Xstory_cloze Lin et al. (2021) for multi-turn dialogues and multilingual tasks. For our GRPO experiments, we focus on mathematical reasoning. We use GSM8K Cobbe et al. (2021) for math reasoning capabilities and Geometry3k Lu et al. (2021) for multimodal mathematical reasoning.
Prompt Sets. As detailed in Section 4.1, we constructed distinct prompt sets for training and evaluation. The training set contains 400 diverse prompts generated exclusively via LLMs; this approach demonstrates that PAFT can fully automate the construction of training materials without manual intervention. For evaluation, we carefully designed a separate test set containing 50 prompts that includes not only LLM-generated prompts but also intentionally human-written instructions. The inclusion of human-written prompts is crucial for comprehensively validating the generalization of our model capabilities and its practical utility in real-world scenarios. This strict separation between a fully synthetic training set and a hybrid test set ensures a rigorous assessment of our method’s effectiveness. Further details on the prompt generation process are provided in Appendix D.
Baseline Comparisons. We establish five baselines to isolate the impact of prompt engineering on fine-tuning performance: the original pre-trained model (Base Model); the model fine-tuning with human-designed prompts (User) following Wei et al. (2024); the model fine-tuning with the highest accuracy of training prompts (TopAccuracy); the model fine-tuning with BATprompt Shi et al. (2024) most robust prompt (BATprompt); and fine-tuning with ZOPO Hu et al. (2024) optimal prompt selection (ZOPO). All models, including baselines, are evaluated on identical test prompts, enabling direct comparison of performance consistency across methods.
Experimental Setup. To comprehensively evaluate the effectiveness of PAFT, our experiments cover two main training paradigms: supervised fine-tuning (SFT) and reinforcement learning fine-tuning (RLFT). For the SFT paradigm, we adopt Low-Rank Adaptation (LoRA Hu et al. (2022)) as a representative method, and the specific experimental parameters are shown in the Appendix B.2; for RLFT, we employ Group Relative Policy Optimization (GRPO Shao et al. (2024)) and the specific experimental parameters are shown in the Appendix B.3. In these settings, we utilize a series of large language models (LLMs), including Llama3-8B, Llama-3.1-8B, Llama-3.2-3B Meta (2024), Qwen2.5-7B, and Qwen2.5-VL-7B Qwen et al. (2025). Detailed correspondence between model datasets and training paradigms is provided in Appendix Table 7. Our implementation is based on the Llama-factory framework, and all evaluations are performed using OpenCompass. All experiments are conducted on NVIDIA A100, V100, 4090, and L40 GPU clusters. For detailed configuration, see the Appendix B.
5.2 Main Results
Prompt Robustness. As demonstrated across Tables 1, Table 2, Figures 4, 5, 6, and 7, PAFT exhibits remarkably low variance across all evaluation tasks, indicating superior prompt robustness. This enhanced stability stems from our dynamic prompt selection strategy (Sec. 4.2), which continuously adjusts prompts during training, compelling the model to learn essential task features rather than overfitting to specific prompt formats. In contrast, baseline approaches face significant limitations: user prompts rely on manual design with inconsistent quality; TopAccuracy and ZOPO tend to overfit to high-performing training prompts with poor generalization; and while BATprompt addresses robustness, it remains less effective than our method. The low variance of PAFT translates to more stable performance and stronger generalization across diverse prompts, enabling development of more user-friendly QA systems, format-independent agent systems, and directly evaluate the true ability of LLMs by better decoupling the ability from the prompting engineering. Notably, PAFT achieves acceptable performance across most prompts, significantly outperforming all baselines (Table 1, Top column) while maintaining high training efficiency (detailed in Appendix C).
| SFT Model | PAFT Model | Improvement | ||||
| Dataset | Min | Con | Min | Con | Min | Con |
| HellaSwag | 87.20 | 61.26 | 91.30 | 84.61 | +4.10 | +23.35 |
| PIQA | 75.16 | 62.13 | 88.72 | 84.12 | +13.56 | +22.99 |
| HumanEval | 45.26 | 12.46 | 52.89 | 50.16 | +7.63 | +37.70 |
| RACE-mid | 72.36 | 50.16 | 85.07 | 83.27 | +12.71 | +33.11 |
| RACE-high | 71.68 | 49.67 | 84.26 | 81.26 | +12.58 | +31.59 |
| GSM8K | 40.36 | 10.26 | 75.13 | 74.13 | +34.77 | +63.87 |
| Geometry3k | 31.26 | 12.37 | 38.90 | 37.12 | +7.64 | +24.75 |
| T-Eval | 42.13 | 19.26 | 61.27 | 59.17 | +19.14 | +39.91 |
To quantify this robustness under more demanding conditions, we introduce two stringent metrics. First, we measure minimum accuracy (Min) on the test set of 50 unseen prompts to evaluate worst-case performance under normal conditions. Second, we assess conditional accuracy (Con) on a challenging set of 10 adversarially crafted prompts containing paraphrases, misspellings, or other modifications to measure resilience against noisy inputs. As shown in Table 4, the PAFT-trained model achieves significantly higher minimum and conditional accuracy across all datasets. This result underscores the ability of PAFT to maintain effective performance even when faced with substantial prompt perturbations and adversarial noise.
SOTA Performance. As demonstrated across our experiments (Tables 1 and 2; Figures 4–7), PAFT consistently achieves SOTA performance by significantly outperforming existing baselines. This superior performance stems directly from our Dynamic Fine-Tuning algorithm (Algorithm 1; Section 4.2), which effectively decouples the underlying fundamental principles of a task from any specific prompt formulation. This decoupling allows the model to focus on learning the essential features of task, rather than becoming entangled in the nuances of a particular prompt. Ultimately, this process enables the model to learn the fundamental principles of downstream tasks instead of merely overfitting to superficial prompt patterns, which is the key reason why PAFT model can achieve its SOTA generalization and performance.
| # and | Hellaswag | PIQA | Winogrande | RACE-mid | RACE-high | Average |
| = 1, = 3 | 93.58 ( 1.47) | 89.33 ( 0.63) | 81.78 ( 1.11) | 86.30 ( 2.73) | 84.35 ( 2.24) | 87.07 ( 1.64) |
| = 2, = 3 | 93.59 ( 1.24) | 88.37 ( 0.49) | 82.09 ( 0.81) | 86.30 ( 2.64) | 84.02 ( 2.24) | 86.87 ( 1.48) |
| = 4, = 3 | 93.83 ( 1.10) | 89.07 ( 0.53) | 81.96 ( 1.15) | 87.26 ( 2.23) | 85.17 ( 1.71) | 87.46 ( 1.34) |
| = 8, = 3 | 93.83 ( 0.70) | 88.99 ( 0.59) | 82.69 ( 0.97) | 86.25 ( 2.75) | 84.36 ( 2.06) | 87.22 ( 1.41) |
| = 1, = 6 | 93.37 ( 1.47) | 88.32 ( 0.68) | 81.05 ( 3.44) | 84.40 ( 2.30) | 83.34 ( 1.66) | 86.10 ( 1.91) |
Inference Efficiency. PAFT enhances inference efficiency not by accelerating per-token generation speed, but by enabling the model to produce correct and concise responses using significantly fewer tokens. Our measurements across all test prompts and datasets (Table 3) demonstrate that this token reduction leads to consistently faster overall inference times compared to baseline methods. This efficiency stems from the prompt robustness of model, which is a core outcome of our training approach. Unlike baseline models that may overfit to superficial prompt patterns, PAFT learns underlying task principles. As a result, even in the face of changing or modified prompts, the model remains focused on the fundamental task objectives, avoiding the need to output redundant dialogue content, such as understanding unknown instructions. This making PAFT particularly valuable for real-world applications requiring rapid responses, such as dialogue systems and AI agents, while simultaneously reducing computational resource requirements. See Appendix C.2 for more detail.
5.3 Ablation Studies
Hyperparameter Robustness. This ablation study demonstrates the robustness of PAFT to the hyperparameters (iterations per prompt) and (epochs). As shown in Table 5, PAFT achieves stable performance across a broad range of (1 to 8) and (3 to 6) values, with minimal fluctuations in accuracy and variance. Notably, PAFT achieves near-optimal performance with default settings (, ), attaining an average accuracy of 87.46%() across all tasks. This robustness reduces the need for extensive hyperparameter tuning, making PAFT a practical and efficient solution for real-world applications.
Impact of Training Prompt Quantity. We conduct an ablation study to investigate the impact of varying numbers of training prompts on model performance, thus validating the effectiveness of PAFT. The experimental results, shown in Figure 8, demonstrate that as the number of prompts increases, the average accuracy of the model significantly improves, while the standard deviation decreases, indicating more stable and reliable performance. However, the performance gains diminish as the number of prompts increases, with only marginal improvements observed beyond a certain threshold. This suggests that while adding prompts can enhance performance, PAFT achieves competitive results with a minimal number of prompts, rendering excessive prompts unnecessary. In most cases, PAFT achieves strong performance with as few as 10 high-quality prompts, and further increases yield only marginal gains. The efficiency of PAFT is particularly notable, as it delivers excellent performance with a minimal number of prompts, making it highly suitable for resource-constrained scenarios where computational efficiency is critical. These findings underscore the practicality and efficiency of PAFT, offering a robust and efficient solution for real-world applications.
6 Theoretical Insights
The capability of PAFT to generalize effectively to unseen prompt formulations can be rigorously understood through the lens of domain adaptation theory Ben-David et al. (2006, 2010). In this theoretical construct, the collection of training prompts along with the task-specific training data delineates the source domain. Besdies, the set of novel test prompts , paired with , represents the target domain. PAFT aims to learn a model , where denotes the hypothesis class, by minimizing the empirical risk computed over instances where each prompt is sampled from .
A foundational result from domain adaptation theory Ben-David et al. (2010) provides an upper bound on the expected risk of on the target prompt distribution with :
Here, is the empirical risk on training prompts. The term signifies model complexity (e.g., related to Rademacher complexity Yin et al. (2020)), which typically diminishes as the number of distinct training prompts increase; this term captures the generalization gap on the source domain. The divergence between the training and test prompt distributions is quantified by . Finally, encapsulates the optimal joint error achievable by a hypothesis in on both domains. The key of PAFT is designed to optimize this bound for improved generalization.
Complexity Control. By employing a substantial number of distinct training prompts , PAFT inherently works to reduce the complexity term . This ensures that the model performance observed on the training prompts becomes a more faithful estimator of its true performance across the entire distribution, fostering more stable learning. This effect is empirically supported by our ablation studies in Section 5.3 (Figure 8), which demonstrate improved stability with more prompts.
Domain Alignment. Minimizing the domain discrepancy term is critically dependent on constructing a diverse and comprehensive set of candidate prompts (see Section 4.1 for details). A more diverse is more likely to effectively cover or closely approximate the diverse and unseen distribution of test prompts . This approximation reduces the divergence between the training and test prompt distributions and enhances the transferability of knowledge learned from to .
Proposition 1 (MMD as an Upper Bound on Discrepancy).
The discrepancy term can be bounded by the Maximum Mean Discrepancy (MMD) Gao et al. (2021a) upper bound:
where MMD is defined as:
The proof of Proposition 1 is provided in the Appendix A. Therefore, we can quantify the upper bound of domain difference using MMD. As illustrated in Figure 9, an increasing number of diverse training prompts cover a wider semantic space, bringing closer to . This proximity reduces the upper bound of the target prompt distribution.
Generalization Guarantee. By minimizing the empirical risk across a sufficiently large and varied corpus of prompts, PAFT encourages the model to internalize the underlying task semantics, rather than merely memorizing superficial prompt structures. This principled approach is key to improving the model performance when confronted with novel and unencountered prompts.
7 Conclusion
PAFT offers a compelling solution for enhancing the prompt robustness of LLMs. By dynamically adjusting prompts during fine-tuning, PAFT significantly improves model prompt robustness and performance across diverse prompt formulations. Notably, PAFT boosts inference speed with maintained training cost. This approach paves the way for more reliable and efficient LLM deployment in real-world applications.
Acknowledgement
This work is supported in part by Shenzhen Science and Technology Program under Grant ZDSYS20220527171400002, the National Natural Science Foundation of China (NSFC) under Grants 62271324, 62231020 and 62371309.
Limitations
In this section, we discuss potential limitations of PAFT and outline promising directions for future research. While PAFT demonstrates significant progress in enhancing the prompt robustness of Large Language Models (LLMs), certain aspects warrant further investigation. A key area for improvement lies in the dynamic prompt selection strategy employed during fine-tuning. Currently, PAFT utilizes a random sampling approach, which, while exposing the model to a diverse range of prompts, may not be the most efficient or effective method. Exploring more sophisticated sampling techniques, such as curriculum learning or importance sampling, could potentially optimize the training process and further enhance robustness. For instance, prioritizing prompts that induce higher loss or those that are more representative of the overall prompt distribution could lead to faster convergence and improved generalization. Furthermore, integrating adversarial learning into the dynamic fine-tuning phase presents a compelling avenue for future work. Generating adversarial prompts on-the-fly, perhaps through gradient-based updates, could further challenge the model and encourage it to learn more robust task representations. This approach could be particularly beneficial in mitigating the impact of maliciously crafted or unexpected prompts. However, the well-known instability of adversarial training remains a significant hurdle. Stabilizing the training process, perhaps through techniques like robust optimization or regularization, is crucial for realizing the full potential of this approach. Investigating different adversarial prompt generation strategies and their impact on model robustness would be a valuable contribution.
Ethics Statement
We have manually reevaluated the dataset we created to ensure it is free of any potential for discrimination, human rights violations, bias, exploitation, and any other ethical concerns.
Appendix A Theoretical Proof
We first state a core assumption regarding the richness of the reproducing kernel Hilbert space (RKHS).
Assumption 1 (Richness of RKHS).
We assume that the RKHS generated by the chosen kernel function is large enough to contain all 1-Lipschitz functions. Formally, we assume that there exists a constant such that any function with a Lipschitz constant at most 1 (i.e., ) is contained in and has a bounded norm, i.e., .
Proof Sketch.
The difference term is the supremum over all 1-Lipschitz functions. According to the assumption 1, we can extend this function space to the RKHS sphere of radius , which contains all 1-Lipschitz functions, and then obtain the definition of MMD by rescaling.
This derivation formally shows that we can use MMD to quantize . ∎
Appendix B Experimental setting
In the main experiment, we compared PAFT with the baseline. The datasets and experimental parameters are as follows:
B.1 Dataset
In this section, we introduce the statistics of the dataset. The statistics of the dataset are shown in Table 6.
| Number of samples | train dataset | validation dataset | test dataset |
| Hellaswag | 39900 | 10000 | 10000 |
| PIQA | 16000 | 2000 | 3000 |
| Winogrande | 40398 | 1267 | 1767 |
| RACE | 87866 | 4887 | 4934 |
| Type | Method | LLM | Dataset |
| Knowledge Understanding | SFT | Llama3-8B | HellaSwag |
| Language Understanding | SFT | Llama3-8B | WinoGrande |
| Math Reasoning Capabilities | GRPO | Qwen2.5-7B | GSM8K |
| Reading Reasoning Capabilities | SFT | Llama3-8B | RACE |
| Grounding and Abstractive Summarization | SFT | Llama3-8B | PIQA |
| Coding Capabilities | SFT | Qwen2.5-7B | HumanEval |
| Tool use | SFT | Llama-3.1-8B | T-Eval |
| Multi-turn dialogues and multilingual tasks | SFT | Llama-3.2-3B | Xstory_cloze |
| Multimodal mathematical reasoning | GRPO | Qwen2.5-VL-7B | Geometry3k |
B.2 Specific SFT experimental parameters
Based on the LLaMA3-8B model configuration, several adjustments were made to optimize model performance. In the baseline model experiment, generation parameters were adjusted to ensure the correct output. In the LoRA experiment, adjustments to the generation parameters were retained, and LoRA-related parameters were adjusted. In the PAFT experiment, the size of the validation set was adjusted to control the time required to search for the optimal layer. For specific experimental parameters, see the table 8.
| Methods | LoRA Target | Max Length | SFT Samples | LR | Training Prompts | Epoch |
| LoRA | q & v Proj | 1024 | 20000 | 0.0001 | 1 | 3 |
| PAFT | q & v Proj | 1024 | 20000 | 0.0001 | 400 | 3 |
B.3 PAFT Integration with GRPO
To demonstrate the versatility of our PAFT framework beyond SFT, we also integrated it with the Reinforcement Learning Fine-Tuning (RLFT) paradigm. We selected Group Relative Policy Optimization (GRPO) as a representative RLFT method and applied our dynamic prompting strategy to its training process for the GSM8K dataset. The setup is as follows:
Candidate Prompt Construction. Following the core principle of PAFT, we first constructed a diverse candidate prompt set for mathematical reasoning. We utilized multiple large language models (LLMs) to generate a wide variety of prompts related to math problems, which formed the prompt training set. To ensure a rigorous evaluation, additional prompts from real dialogues and other synthetic sources were used to create a distinct test set, guaranteeing that the prompt training and test sets were entirely distinct.
Dynamic GRPO Training Process. We integrated our dynamic sampling mechanism directly into the GRPO training loop. In each step of the standard GRPO process, instead of using a fixed instruction, we randomly sampled a prompt from our candidate training set. This sampled prompt (e.g., “Please help me solve this math problem {GSM8K_problem}”) was then combined with a problem instance from the GSM8K dataset to form the final input for the model’s computation step. This process aligns with the standard procedure for Soft Fine-Tuning.
Hyperparameters. The key hyperparameters used for our GRPO experiments are detailed in Table 9.
| Parameter | Value |
| Learning Rate | 5e-6 |
| Num Generations | 16 |
| Epochs | 10 |
Appendix C Training cost and inference time
C.1 Training cost
PAFT Maintains Training Efficiency. We now turn our attention to the training efficiency of PAFT. A critical consideration for any practical fine-tuning approach is its impact on training time. Introducing complex mechanisms or additional computational overhead can significantly hinder the training process, especially when dealing with large language models and extensive datasets. Therefore, it is essential to demonstrate that PAFT does not introduce such burdens.
To rigorously evaluate the training time implications of PAFT, we conducted a series of experiments, using Low-Rank Adaptation (LoRA) Hu et al. (2022) as a representative example of a parameter-efficient fine-tuning method. LoRA has gained popularity due to its ability to adapt pre-trained models with minimal computational cost, making it a suitable baseline for our analysis. Our experiments, the results of which are presented in Table 10, directly compare the training time required for traditional LoRA fine-tuning with the training time required for PAFT integrated with LoRA.
The key finding from our analysis is that PAFT does not introduce any noticeable increase in training time. The data in Table 10 clearly demonstrates that the training duration remains virtually identical whether we employ standard LoRA or incorporate PAFT’s dynamic prompt selection mechanism. This crucial observation underscores the efficiency of PAFT. The dynamic prompt selection process, which is central to PAFT’s ability to enhance prompt robustness, is implemented in a way that does not add significant computational overhead. This is because the selection process is lightweight and seamlessly integrated into the existing training loop. Rather than requiring complex computations or extensive data manipulations, PAFT efficiently chooses from a diverse set of prompts, allowing the model to experience a wider range of input formulations without incurring a substantial time penalty. This efficient dynamic prompt selection is critical for the practical applicability of PAFT, ensuring that it can be readily deployed without compromising training efficiency. Furthermore, this efficiency allows for more extensive experimentation and exploration of different prompt variations, ultimately leading to more robust and generalizable models.
| Training time/h | Hellaswag | PIQA | Winogrande | RACE | Average |
| LoRA + user-specified prompt | 3.01 | 2.35 | 3.27 | 3.95 | 3.15 |
| LoRA + TopAccuracy prompt | 3.00 | 2.29 | 2.98 | 3.93 | 3.05 |
| LoRA + BATprompt | 3.02 | 2.23 | 3 | 3.93 | 3.05 |
| LoRA + ZOPO prompt | 2.97 | 2.3 | 2.97 | 3.83 | 3.02 |
| PAFT | 2.98 | 2.32 | 3.38 | 3.81 | 3.12 |
Efficient Candidate Prompt Generation. A key aspect of PAFT’s effectiveness lies in its ability to generate a diverse and high-quality set of candidate prompts efficiently. The process of constructing these candidate prompts involves leveraging the capabilities of external large language models (LLMs), which naturally raises the question of associated costs. Specifically, we sought to quantify the token usage required for candidate prompt generation, as this directly translates to the expense incurred when interacting with commercial LLM APIs.
To address this, we conducted a detailed analysis of the token consumption during the candidate prompt generation phase of PAFT. Our investigation, the results of which are summarized in Table 11, focuses on the number of tokens required to produce a sufficient variety of prompts suitable for subsequent selection and fine-tuning. We meticulously tracked the token usage across various prompts generated for different tasks, considering factors such as prompt length, complexity, and diversity.
The findings presented in Table 11 demonstrate that PAFT requires remarkably few tokens to generate a substantial pool of candidate prompts. This efficiency stems from PAFT’s strategic approach to prompt engineering. Rather than relying on brute-force generation or computationally intensive search methods, PAFT employs a carefully designed prompting strategy that encourages the external LLMs to produce a wide range of prompt formulations with minimal token consumption. This is achieved through techniques such as few-shot prompting with carefully chosen examples, targeted instructions that guide the LLM towards desired prompt characteristics, and potentially iterative refinement of prompts based on preliminary evaluation. The low token count is crucial for practical applications, as it minimizes the cost associated with using commercial LLM APIs. Moreover, this efficiency enables the exploration of a broader range of potential prompts within a fixed budget, increasing the likelihood of discovering highly effective prompts that contribute to improved model robustness. This efficient prompt generation process is a significant advantage of PAFT, enabling it to achieve superior performance without incurring prohibitive costs.
| Tokens | Hellaswag | PIQA | Winogrande | RACE | Average |
| Total Tokens | 11.7k | 12.1k | 10.9k | 12.3k | 11.75k |
C.2 Inference Times
In this section, we show some special cases to demonstrate that PAFT improves reasoning efficiency by improving the model’s understanding of the core task semantics, enabling concise and accurate responses with fewer tokens.
Appendix D Prompt
In this section, we present a selection of training and test prompts to illustrate the efficacy of our prompt construction algorithm and to provide a clearer understanding of operational process of PAFT. Due to space constraints, we only list 10 prompts as examples. Section D.1 shows how we guide LLMs to generate candidate prompts. Section D.2 showcases examples of training prompts, Section D.3 highlights test prompts, and Section D.4 outlines the prompts utilized by the baseline method.
D.1 Automated Prompt Generation Strategy
To construct a diverse and high-quality set of candidate prompts, we employ a strategy that leverages large language models (LLMs) through two distinct templating approaches: zero-shot and one-shot prompting. These templates are designed to be general, requiring only minor modifications to synthesize prompt sets for various datasets.
Zero-shot Prompting. Our zero-shot approach uses a general template that instructs an LLM to generate multiple prompt variations for a given task type, without being constrained by a specific problem instance. This method is effective for tasks where the output format is straightforward. For example, to generate prompts for a commonsense reasoning task like PIQA, we use the following instruction:
One-shot Prompting. For tasks requiring a specific output format, such as the step-by-step reasoning in mathematical problems, we use a one-shot template. This template provides the LLM with an explicit example of the desired output structure in addition to the generation instructions, thereby guiding the model’s response format[cite: 1]. For datasets like GSM8K and Geometry3K, our one-shot instruction is as follows:
This templating strategy is designed to produce prompts that are both adaptable and general. By crafting instructions that elicit the necessary information without being overly task-specific, we ensure the generated prompts can be applied across different datasets with minimal modification. This approach is fundamental to our goal of enhancing prompt robustness and practical applicability, demonstrating that our framework can automate the creation of effective and varied training prompts.
D.2 Train prompt
In this section, we present the prompts generated using the method outlined in Section 4.1 across various datasets. All prompts listed here are utilized for training purposes.
D.3 Test prompt
In this section, we present the prompts generated using the method outlined in Section 4.1 across various datasets. All prompts listed here are utilized for testing purposes, and they are not visible during training.
D.4 Baseline prompt
In this section, we present the best prompts generated or filtered using the baseline for training.